If a bullet is fired with a speed of 50m/s at a 45° angle, what is the height of the bullet when its direction of motion becomes a 30° angle with the horizontal plane?

Grade:11

1 Answers

Arun
6 years ago
Dear Aditi
 
So we know that the total initial speed is 50m/s then the horizontal component will have to be 50m/s * cos (45) = 35.36 m/s. Now to find the vertical component that would make a 30 degree angle with 35.36 m/s as the horizontal component we use the formula tan (30) = y/35.56 then solving for y → y = tan (30) 35.56 = 20.41 m/s. Then we use the pythagorean theorem to combine the components sqrt (35.56^2 + 20.41^2) = 41.00 m/s.
 
Regards
Arun (askIITians forum expert)